OpenStudy (anonymous):

as nin types out his humongous reply, try this integral\[\int \limits_0^2\sqrt{x+ \sqrt{x+\sqrt{x+\cdots}}}dx\]

OpenStudy (anonymous):

if we call this expression \(y\) then \(y^2 - y = x\) so probably a substitution like \(t^2 - t = x\) should work? i'm not sure

ganeshie8 (ganeshie8):

I'm getting \(\dfrac{19}{6}\)

OpenStudy (anonymous):

oh great, how? i'm just not good at integrals in general.

ganeshie8 (ganeshie8):

I have used your method and solved \(y\)

OpenStudy (anonymous):

quadratic equation?

OpenStudy (ikram002p):

And dx with respect to y would be?

ganeshie8 (ganeshie8):

Yes \[\sqrt{x+ \sqrt{x+\sqrt{x+\cdots}}}=y \\~\\\implies x+y = y^2\\~\\\implies y = \frac{1}{2}(1+\sqrt{4x+1}) \]

OpenStudy (anonymous):

simple... i'll post more

hartnn (hartnn):

\(1- \sqrt{4x+1}\) can never be positive or 0 ? why wasn't it considered....

hartnn (hartnn):

i know y must be positive..

ganeshie8 (ganeshie8):

when x = 0, the integral evaluates to 0. for x > 0, the value of integrand is clearly positive. \(1-\sqrt{4x+1}\)is never positive in the given range. so yeah we simply reject this root

hartnn (hartnn):

gotcha :)

OpenStudy (anonymous):

OH MY GOD I JUST REALISED I DERIVED AN EULER SUB ABOVE. LOL

OpenStudy (anonymous):

"if we call this expression \(y\) then \(y^2 - y = x\) so probably a substitution like \(t^2 - t = x\) should work? i'm not sure" that is an euler's substitution